# General spin models from noncollinear spin density functional theory and spin-cluster expansion

### v1

@@ Abstract
We present a data-efficient framework for constructing general classical spin Hamiltonians from the spin-cluster expansion (SCE) combined with fully self-consistent noncollinear spin density functional theory (DFT). The key idea is to fit an SCE model to magnetic torques rather than to total energies. Because torques are site-resolved vectors, each configuration supplies many independent constraints, which makes the regression well conditioned and sharply reduces the number of DFT calculations needed, especially in large supercells. Applied to the B20-type chiral magnets ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$, the resulting models nonperturbatively extract the full pairwise exchange tensor (isotropic exchange, anisotropic symmetric exchange, and the Dzyaloshinskii–Moriya interaction) and predict helical spin period via a micromagnetic mapping. The composition trends and the divergence of the period near the chirality sign change are reproduced in line with experiments. Because the SCE framework is systematic, it also enables systematic analysis of interaction order; training on increasingly disordered spin configurations shows that the lowest-order model loses torque accuracy, whereas including higher-order interactions restores predictive power. These advances enable near-DFT-accurate spin models for finite-temperature magnetism and complex textures at modest data cost, while providing a systematic, extensible, and nonperturbative route to quantitative first-principles parameterization and predictive materials design. An open-source implementation is available as the Julia package, Magesty.jl.
@@ I Introduction
Recent advances in spintronic materials and technologies have been substantial, laying the groundwork for ultrafast, energy-efficient devices that exploit atomic-scale magnetic order and spin dynamics [1, 2]. In parallel, first-principles methods—especially density functional theory (DFT) [3, 4]—have become central to the understanding and design of magnetic and spintronic materials. However, fully ab initio treatments often reach practical limits when addressing finite-temperature magnetism and mesoscopic spin dynamics. Consequently, a broad class of statistical-mechanical models that coarse-grain the spin degrees of freedom has been developed [5, 6, 7, 8, 9, 10]. A natural coarse-graining is to map the electronic system onto a classical spin model—a classical spin Hamiltonian that retains the orientations of local magnetic moments as degrees of freedom while integrating out charge and longitudinal fluctuations—so that the energy of a spin configuration is expressed through effective interaction parameters.
To leading order, the spin Hamiltonian in the vicinity of a locally stable spin configuration takes the form
$E=\sum_{i,j,i\neq j}\sum_{\alpha,\beta}J_{ij}^{\alpha\beta}\,e_{i\alpha}e_{j\beta},$ (1)
where the Greek indices $\alpha,\beta$ denote Cartesian components, $e_{i\alpha}$ is the $\alpha$ component of the unit vector $\hat{\bm{e}}_{i}$ specifying the moment direction at site $i$, and $J_{ij}^{\alpha\beta}$ is the two-body exchange interaction tensor between sites $i$ and $j$. This form is commonly sufficient to describe the immediate neighborhood of a locally stable magnetic state. However, when one seeks a unified description over a broader set of spin configurations, higher-order terms become necessary. For example, temperature-dependent effective two-body isotropic terms (the Heisenberg interactions)—arising from the renormalization of higher-order interactions—are essential for modeling finite-temperature magnetic energetics [11]. Furthermore, recent studies suggest that higher-order interactions have an important role for the stabilization of complex spin structures, such as chiral spin textures and skyrmions [12, 13, 14]. Accordingly, the importance of establishing quantitative methods to evaluate higher-order interactions is increasingly pronounced.
@@
with $n_{\mathcal{C}}$ the number of sites in cluster $\mathcal{C}$ and ${\bm{l}}=(l_{1},\ldots,l_{n_{\mathcal{C}}})$, ${\bm{m}}=(m_{1},\ldots,m_{n_{\mathcal{C}}})$. In spherical coordinates $(\theta,\phi)$,
@@
The set $\Phi_{\mathcal{C}{\bm{l}}{\bm{m}}}$}, together with the constant term, forms an orthogonal and complete basis set [15]. By increasing the maximum degree $l$ and the cluster size, the model can be systematically improved.
@@
In principle, the SCE coefficients $J_{\mathcal{C}{\bm{l}}{\bm{m}}}$ can be obtained by regressing to constrained noncollinear-spin DFT total energies, as shown in earlier SCE work [17] and by analogy with the cluster expansion for substitutional alloys [18]. In practice, however, magnetic interactions operate on energy scales far smaller than chemical interactions in alloys, which demands higher precision from DFT. Moreover, whereas alloy variables are discrete, spin orientations are continuous, so the configuration space is vast; fits that rely only on total energies therefore require prohibitively many samples to reproduce the energy landscape with accuracy. A practical solution is to augment the regression with information beyond total energies—namely, derivatives with respect to spin orientations (magnetic torques). The use of derivatives is well established elsewhere: in finite-displacement phonon calculations one fits DFT-computed forces to build Taylor-expanded potentials [19, 20], and modern machine-learning interatomic potentials train on both energies and forces [21]. In magnetism, analogous strategies—fitting to magnetic torques from noncollinear-spin DFT—have been applied primarily to some spin models [22, 23, 24, 25]. Integrating these ideas into the SCE framework yields a data-efficient and systematic route derivation of general spin models.
@@
In this work, we present an efficient and scalable spin-cluster expansion (SCE) framework fitted to magnetic torque data. Unlike earlier SCE implementations [26, 17, 27], our approach is able to treat spin–orbit coupling (SOC) explicitly, which breaks global spin rotational invariance and necessitates anisotropic terms in spin models. This enables nonperturbative evaluation of arbitrary anisotropic terms in the spin Hamiltonian, including single-ion anisotropy, symmetric exchange anisotropy, and the antisymmetric Dzyaloshinskii–Moriya interaction (DMI). The paper is organized as follows. Section II.1 introduces a formulation in real spherical harmonics that ensures a real-valued energy and streamlines SOC treatment. Building on this, Section II.2 derives symmetry-adapted basis functions via a projection operator to enforce crystal-symmetry constraints on SCE coefficients. Section II.3 then describes the torque-fitted regression scheme, and Section II.4 presents the real harmonics and their derivatives on the unit sphere in Cartesian coordinates used in the torque-fitted regression. In Sec. II.5, we describe a simple mean-field-based sampling scheme that generates temperature-controlled spin configurations to serve as training data for constructing the SCE model. We apply our method to B20-type chiral compounds ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$ (Section II.6) as demonstrations. First, we show that the torque-fitted regression drastically reduces the number of DFT data points needed to construct an SCE model. We then use the resulting models to evaluate the full exchange tensor (consisting of isotropic exchange, anisotropic symmetric exchange, and DMI), extract the helical period via micromagnetic models, and reproduce the experimentally observed divergence of the period as a function of composition. Finally, by increasing the sampling temperature to generate more strongly fluctuating spin configurations, we assess the validity of a lowest-order spin Hamiltonian and observe reduced predictive accuracy for torque predictions, indicating that higher-order interactions are required at elevated temperatures. The paper concludes with Section III, which summarizes the main findings and outlines future directions. Our implementation is released as an open-source Julia package, Magesty.jl.
II Results
@@ II.1 Spin-cluster expansion with real spherical harmonics
In the original implementation of SCE [15, 16], complex spherical harmonics were employed. Subsequently, Ref. [26] ensured a real-valued energy by assuming global spin rotation invariance (i.e., neglecting SOC) together with time-reversal symmetry. When SOC is included, however, using complex spherical harmonics requires explicit constraints on the SCE coefficients to guarantee that the total energy remains real. To avoid this bookkeeping and to simplify SOC implementations, we work with real spherical harmonics.
@@
where $N$ is the total number of sites and $\int\mathop{}\!\mathrm{d}\hat{\bm{e}}_{i}$ denotes integration over the unit sphere [15, 26]. One then obtains the orthonormality relation
[…]
@@ II.2 Construction of symmetry-adapted basis functions
We consider a paramagnetic group $\bar{G}_{0}$ which consists of a space group $G_{0}$ and the time-reversal operation $\hat{\mathcal{T}}$ as
@@
where $\hat{\mathcal{T}}G_{0}$ is the set formed by the product of the time-reversal operation $\hat{\mathcal{T}}$ and the symmetry operations belonging to the space group $G_{0}$. Note that we do not impose any symmetry of a specific magnetic order; magnetic space-group symmetries are not considered. Magnetic space groups may emerge when the original spin Hamiltonian is approximated by an effective field Hamiltonian (e.g., by mean-field approximation), but they do not represent the fundamental symmetry group that the original spin model is required to obey. The projection operator of the identity irreducible representation $\hat{P}_{\rm id}$ is defined as follows:
@@
It follows that only cluster terms with an even $\sum_{i=1}^{n_{\mathcal{C}}}l_{i}$ are $\hat{\mathcal{T}}-$even and therefore allowed as a time-reversal-invariant function, whereas terms with an odd $\sum_{i=1}^{n_{\mathcal{C}}}l_{i}$ are $\hat{\mathcal{T}}-$odd and thus excluded [26]. Combining this with the spatial symmetry operations, we can construct the projector onto the identity irreducible representation. The symmetry-adapted basis functions are the eigenfunctions of $\hat{P}_{\rm id}$ with eigenvalue $1$ [28].
@@ II.3 Estimation of SCE coefficients utilizing local magnetic torques
$\bm{T}_{{\rm DFT},i}\equiv-\hat{\bm{e}}_{i}\times\!\left(-\,\nabla_{\hat{\bm{e}}_{i}}E_{\rm DFT}\right)\;=\;-\bm{M}_{i}\times\bm{B}_{i},$ (25)
where $E_{\rm DFT}$ is the total energy, $\bm{M}_{i}=M_{i}\hat{\bm{e}}_{i}$ is the local magnetic moment, and $\bm{B}_{i}$ is a site-averaged effective magnetic field. At self-consistency, $\bm{B}_{i}$ is transverse to $\hat{\bm{e}}_{i}$ and equal in magnitude and opposite in direction to the constraining field [29]. Hence both sides of $\nabla_{\hat{\bm{e}}_{i}}E_{\rm DFT}=M_{i}\bm{B}_{i}$ lie in the tangent plane at $\hat{\bm{e}}_{i}$. For completeness, the gradient on the unit sphere with respect to $\hat{\bm{e}}_{i}=(e_{ix},e_{iy},e_{iz})$ is
@@
$\bm{T}_{{\rm SCE},i}=-\hat{\bm{e}}_{i}\times\bigl(-\nabla_{\hat{\bm{e}}_{i}}E_{\rm SCE}\bigr)=-\hat{\bm{e}}_{i}\times\Bigl[-\bigl(\nabla_{\hat{\bm{e}}_{i}}\bm{u}^{\mathsf{T}}\bigr)\,\bm{J}\Bigr].$ (29)
@@
$\mathbf{A}\;=\;\begin{pmatrix}\hat{\bm{e}}_{1}\times\bigl(\nabla_{\hat{\bm{e}}_{1}}u_{1}\bigr)&\hat{\bm{e}}_{1}\times\bigl(\nabla_{\hat{\bm{e}}_{1}}u_{2}\bigr)&\cdots&\hat{\bm{e}}_{1}\times\bigl(\nabla_{\hat{\bm{e}}_{1}}u_{N_{\rm I}}\bigr)\\ \hat{\bm{e}}_{2}\times\bigl(\nabla_{\hat{\bm{e}}_{2}}u_{1}\bigr)&\hat{\bm{e}}_{2}\times\bigl(\nabla_{\hat{\bm{e}}_{2}}u_{2}\bigr)&\cdots&\hat{\bm{e}}_{2}\times\bigl(\nabla_{\hat{\bm{e}}_{2}}u_{N_{\rm I}}\bigr)\\ \vdots&\vdots&\ddots&\vdots\\ \hat{\bm{e}}_{N_{\rm a}}\times\bigl(\nabla_{\hat{\bm{e}}_{N_{\rm a}}}u_{1}\bigr)&\hat{\bm{e}}_{N_{\rm a}}\times\bigl(\nabla_{\hat{\bm{e}}_{N_{\rm a}}}u_{2}\bigr)&\cdots&\hat{\bm{e}}_{N_{\rm a}}\times\bigl(\nabla_{\hat{\bm{e}}_{N_{\rm a}}}u_{N_{\rm I}}\bigr)\end{pmatrix}\!$.
@@
When the system is well determined, $\bm{J}$} is determined as ${\bm{J}}=(\widetilde{\bf A}^{\mathsf{T}}\widetilde{\bf A})^{-1}\widetilde{\bf A}^{\mathsf{T}}\widetilde{\bm{T}}_{\rm DFT}$.

### v2

@@ Abstract
We present a data-efficient framework for constructing general classical spin Hamiltonians by combining the spin-cluster expansion (SCE) with fully self-consistent noncollinear spin density functional theory (DFT). The key idea is to fit the SCE model to magnetic torques rather than to total energies. Because torques are site-resolved vectors, each spin configuration provides many informative regression targets, improving conditioning and substantially reducing the number of required DFT calculations, especially for large supercells. Applied to the B20-type chiral magnets ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$, the resulting SCE models determine full pairwise exchange tensors—including isotropic exchange, symmetric anisotropic exchange, and the Dzyaloshinskii–Moriya interaction—and predict the helical spin period via a micromagnetic mapping. The composition trends and the divergence of the period at the chirality sign-change point are well reproduced, in agreement with experiment. Moreover, the systematic nature of SCE enables controlled assessment of interaction order: as the training spin configurations become more disordered, the lowest-order model loses torque accuracy, whereas including higher-order interactions restores predictive power. These advances enable near-DFT-accurate spin models for finite-temperature magnetism and complex spin textures at modest computational cost, providing an extensible route to quantitative first-principles parameterization and predictive materials design. An open-source implementation is available as a Julia package, Magesty.jl.
@@ I Introduction
Recent advances in spintronic materials and technologies have been substantial, laying the groundwork for ultrafast, energy-efficient devices that exploit atomic-scale magnetic order and spin dynamics [1, 2]. In parallel, first-principles methods—especially density functional theory (DFT) [3, 4]—have become central for the understanding and design of magnetic and spintronic materials. However, fully ab initio treatments often reach practical limits when addressing finite-temperature magnetism and mesoscopic spin dynamics. Consequently, a broad class of statistical-mechanical models that coarse-grain the spin degrees of freedom has been developed [5, 6, 7, 8, 9, 10, 11]. A natural coarse-graining is to map the electronic system onto a classical spin Hamiltonian that retains only the orientations of local magnetic moments as degrees of freedom. The remaining electronic degrees of freedom, including charge and longitudinal-amplitude fluctuations, are thereby absorbed into effective interaction parameters that reproduce the energy of a given spin configuration.
In the immediate neighborhood of a locally stable magnetic configuration, a pairwise exchange-tensor model is often sufficient:
$E=\sum_{i,j}\sum_{\alpha,\beta}J_{ij}^{\alpha\beta}\,e_{i\alpha}e_{j\beta},$ (1)
where the Greek indices $\alpha,\beta$ denote Cartesian components, $e_{i\alpha}$ is the $\alpha$ component of the unit vector $\hat{\bm{e}}_{i}$ specifying the moment direction at site $i$, and $J_{ij}^{\alpha\beta}$ is the two-body exchange interaction tensor between sites $i$ and $j$. Such a lowest-order description, however, can become insufficient when one seeks a unified description over a broader set of spin configurations, in which case higher-order terms become necessary. For example, temperature-dependent effective two-body isotropic terms (the Heisenberg interactions)—arising from the renormalization of higher-order interactions—can be important for modeling finite-temperature magnetic energetics [12]. Furthermore, recent studies suggest that higher-order interactions have an important role for the stabilization of complex spin structures, such as chiral spin textures and skyrmions [13, 14, 15]. Accordingly, the importance of establishing quantitative methods to evaluate higher-order interactions is increasingly pronounced.
@@
with $n_{\mathcal{C}}$ the number of sites in cluster $\mathcal{C}$ and ${\bm{l}}=(l_{1},\ldots,l_{n_{\mathcal{C}}})$, ${\bm{m}}=(m_{1},\ldots,m_{n_{\mathcal{C}}})$. Here $\bm{l}>0$} is understood componentwise, i.e., $l_{i}\geq 1$ for all sites in the cluster (not merely ${\bm{l}}\neq{\bm{0}}$). In spherical coordinates $(\theta,\phi)$,
@@
The set $\{{\Phi_{\mathcal{C}{\bm{l}}{\bm{m}}}}\}$, together with the constant term, forms an complete orthonormal basis set [16]. By increasing the maximum degree $l$ and the cluster size, the model can be systematically refined to better reproduce the underlying energy landscape.
@@
In principle, the SCE coefficients $J_{\mathcal{C}{\bm{l}}{\bm{m}}}$ can be obtained by regressing to constrained noncollinear-spin DFT total energies, as shown in earlier SCE work [18] and by analogy with the cluster expansion for substitutional alloys [19]. In practice, however, magnetic interactions operate on energy scales far smaller than chemical interactions in alloys, which demands higher precision from DFT. Moreover, whereas alloy variables are discrete, spin orientations are continuous, so the configuration space is vast; fits that rely only on total energies therefore require prohibitively many samples to reproduce the energy landscape with accuracy. A practical solution is to augment the regression with information beyond total energies—namely, derivatives of the total energy with respect to spin orientations (magnetic torques). The use of derivatives is well established elsewhere: in finite-displacement phonon calculations one fits DFT-computed atomic forces to build Taylor-expanded potentials [20, 21], and machine-learning interatomic potentials train on both energies and forces [22]. In magnetism, analogous strategies—fitting to magnetic torques from noncollinear-spin DFT—have been applied primarily to some spin models [23, 24, 25, 11]. Integrating these ideas into the SCE framework yields a data-efficient and systematic route to the derivation of general spin models.
@@
In this work, we present an efficient and scalable SCE framework fitting to magnetic torque data. Unlike earlier SCE implementations [26, 18, 27], our approach is able to treat spin–orbit coupling (SOC) explicitly, which breaks global spin rotational invariance and necessitates anisotropic terms in spin models. This enables nonperturbative evaluation of arbitrary anisotropic terms in the spin Hamiltonian, including single-ion anisotropy, symmetric exchange anisotropy, and the antisymmetric Dzyaloshinskii–Moriya interaction (DMI). The paper is organized as follows. Section II presents the theoretical framework: we formulate SCE using real spherical harmonics to ensure a real-valued energy in the presence of SOC (Sec. II.1), construct symmetry-adapted basis functions to enforce crystal symmetries (Sec. II.2), and describe the torque-fitting regression together with the Cartesian-polynomial representation of real spherical harmonics used to evaluate torque derivatives (Secs. II.3 and II.4). Section III applies the method to B20-type chiral magnets ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$. After summarizing computational details (Sec. III.1) and our mean-field-based sampling scheme for generating temperature-controlled training spin configurations (Sec. III.2), we demonstrate the data efficiency of torque fitting (Sec. III.4), derive micromagnetic parameters and the helical period (Sec. III.5), and assess the role of higher-order interactions using more strongly disordered configurations (Sec. III.6). Finally, Sec. IV summarizes the main findings and discusses possible extensions. Our implementation is released as an open-source Julia package, Magesty.jl.
II Theory
@@ II.1 Spin-cluster expansion with real spherical harmonics
In the original implementation of SCE [16, 17], complex spherical harmonics were employed. Subsequently, Ref. [26] ensured a real-valued energy by assuming global spin rotation invariance (i.e., neglecting SOC) together with time-reversal symmetry. When SOC is included, however, using complex spherical harmonics requires explicit constraints on the SCE coefficients to guarantee that the total energy remains real. To avoid this bookkeeping and to simplify implementations, we work with real spherical harmonics.
@@
where $N$ is the total number of sites and $\int\mathop{}\!\mathrm{d}\hat{\bm{e}}_{i}$ denotes integration over the unit sphere [16, 26].
@@ II.2 Construction of symmetry-adapted basis functions
We consider a paramagnetic (gray) group $\bar{G}_{0}$ which consists of a space group $G_{0}$ and the time-reversal operation $\hat{\mathcal{T}}$ as
@@
where $\hat{\mathcal{T}}G_{0}$ is the set formed by the product of the time-reversal operation $\hat{\mathcal{T}}$ and the symmetry operations belonging to the space group $G_{0}$. Note that we do not impose any symmetry of a specific magnetic order; magnetic space-group symmetries are not considered. Magnetic space groups may emerge when the original spin Hamiltonian is approximated by an effective field Hamiltonian (e.g., by the mean-field approximation), but they do not represent the fundamental symmetry group that the original spin model is required to obey. The projection operator of the identity irreducible representation $\hat{P}_{\rm id}$ is defined as follows:
@@
It follows that only cluster terms with an even $\sum_{i=1}^{n_{\mathcal{C}}}l_{i}$ are $\hat{\mathcal{T}}-$even and therefore allowed as a time-reversal-invariant function, whereas terms with an odd $\sum_{i=1}^{n_{\mathcal{C}}}l_{i}$ are $\hat{\mathcal{T}}-$odd and thus excluded [26]. Combining this with the spatial symmetry operations, we can construct the projector onto the identity irreducible representation. The symmetry-adapted basis functions are the eigenfunctions of $\hat{P}_{\rm id}$ with eigenvalue $1$ [28]. In the actual calculations, we used Spglib library for the evaluation of crystal symmetry [29].
@@ II.3 Estimation of SCE coefficients utilizing local magnetic torques
$\bm{T}_{{\rm DFT},i}\equiv\hat{\bm{e}}_{i}\times\!\left(-\,\nabla_{\hat{\bm{e}}_{i}}E_{\rm DFT}\right)\;=\;\bm{\mathfrak{m}}_{i}\times\bm{B}_{i},$ (24)
where $E_{\rm DFT}$ is the total energy, ${\mathfrak{m}}_{i}=|{\mathfrak{m}}_{i}|\hat{\bm{e}}_{i}$ is the local magnetic moment, and $\bm{B}_{i}$ is a site-averaged effective magnetic field. At self-consistency, $\bm{B}_{i}$ is transverse to $\hat{\bm{e}}_{i}$ and equal in magnitude and opposite in direction to the constraining field [30]. Hence both sides of $\nabla_{\hat{\bm{e}}_{i}}E_{\rm DFT}=|{\mathfrak{m}}_{i}|\bm{B}_{i}$ lie in the tangent plane at $\hat{\bm{e}}_{i}$ [24]. For completeness, the gradient on the unit sphere with respect to $\hat{\bm{e}}_{i}=(e_{ix},e_{iy},e_{iz})$ is
@@
$\bm{T}_{{\rm SCE},i}=\hat{\bm{e}}_{i}\times\bigl(-\nabla_{\hat{\bm{e}}_{i}}E_{\rm SCE}\bigr)=\hat{\bm{e}}_{i}\times\Bigl[-\bigl(\nabla_{\hat{\bm{e}}_{i}}\bm{u}^{\mathsf{T}}\bigr)\,\bm{J}\Bigr].$ (28)
@@
$\mathbf{A}\;=\;\begin{pmatrix}-\hat{\bm{e}}_{1}\times\bigl(\nabla_{\hat{\bm{e}}_{1}}u_{1}\bigr)&-\hat{\bm{e}}_{1}\times\bigl(\nabla_{\hat{\bm{e}}_{1}}u_{2}\bigr)&\cdots&-\hat{\bm{e}}_{1}\times\bigl(\nabla_{\hat{\bm{e}}_{1}}u_{N_{\rm I}}\bigr)\\ -\hat{\bm{e}}_{2}\times\bigl(\nabla_{\hat{\bm{e}}_{2}}u_{1}\bigr)&-\hat{\bm{e}}_{2}\times\bigl(\nabla_{\hat{\bm{e}}_{2}}u_{2}\bigr)&\cdots&-\hat{\bm{e}}_{2}\times\bigl(\nabla_{\hat{\bm{e}}_{2}}u_{N_{\rm I}}\bigr)\\ \vdots&\vdots&\ddots&\vdots\\ -\hat{\bm{e}}_{N_{\rm a}}\times\bigl(\nabla_{\hat{\bm{e}}_{N_{\rm a}}}u_{1}\bigr)&-\hat{\bm{e}}_{N_{\rm a}}\times\bigl(\nabla_{\hat{\bm{e}}_{N_{\rm a}}}u_{2}\bigr)&\cdots&-\hat{\bm{e}}_{N_{\rm a}}\times\bigl(\nabla_{\hat{\bm{e}}_{N_{\rm a}}}u_{N_{\rm I}}\bigr)\end{pmatrix}\!$.
@@
When the system is well determined, the least-squares estimate of $\bm{J}$ is given by ${\bm{J}}=(\widetilde{\bf A}^{\mathsf{T}}\widetilde{\bf A})^{-1}\widetilde{\bf A}^{\mathsf{T}}\widetilde{\bm{T}}_{\rm DFT}$.